OpenStudy (anonymous):

Given: Angle ABD = 10 degrees, Angle ABC = 100 degrees, Angle EFY = 70 degrees 20', ANGLE XFY = 19 degrees 40' Prove Angle DBC is congruent to angle XFE

OpenStudy (anonymous):

@DebbieG 

OpenStudy (anonymous):

And if you can't help it's ok :)

OpenStudy (debbieg):

OK, I think this one just really is about adding/subtracing angle measures, some of which involve partial degrees, e.g., minutes, right? You need to show that the measure of XFE=DBC XFE=XFY+EFY right? So add those: 70*20'+19*40' = ? DBC=ABC-ABD right? So subtract: 100* - 10* =? That should about do it. :)

OpenStudy (anonymous):

Yes So 70*20' + 19*40' = 90* 100*-10*=90 Is that right?

OpenStudy (debbieg):

Yeppers! :)

OpenStudy (debbieg):

Because there are 60min in 1deg.... so the 70*20'+19*40' = (70+19)* + (20+40)'=89*+60'=89*+1*=90*

OpenStudy (anonymous):

Ok! Thank you soo much :) You're a big help

OpenStudy (debbieg):

You're welcome, happy to help! :)

OpenStudy (anonymous):

Do you think it's ok if once in a while I tag you in some problems? I think you're really good at helping and explaining problems

OpenStudy (debbieg):

Aww, thank you! Sure, feel free! Just know that sometimes I am logged in but nowhere near my computer, lol. :)